OpenStudy (anonymous):

Which number would you add to one and nine-tenths to give an approximate sum of 5? three and two-thirds one and sielevenths four and one-twelfth two and seven-eighths

OpenStudy (anonymous):

can sme one help me

OpenStudy (anonymous):

3 2/3 + 9/11 = ? 1 6/11 + 9/11 =? 4 1/12 + 9/11 =? 2 7/8 + 9/11 = ? Which option gives you answer closer to 5?

OpenStudy (anonymous):

ithink A or 8

OpenStudy (anonymous):

A would be 3.67+0.82 = 4.49 B would be 1.55 + 0.82 = 2.37 C would be 4.08 + 0.82 = 4.90 D would be 2.88 + 0.82 = 3.70 Which option is closest to 5?

OpenStudy (anonymous):

c